Assume a research firm collects survey sales data that reveal the relationship between the possible selling prices of hamburgers and the quantity of hamburgers consumers would purchase per year at alternative prices. The report states that if the price of a hamburger is $4, 20,000 will be bought. However, at a price of $3, 40,000 hamburgers will be bought. At $2, 60,000 hamburgers will be bought, and at $1, 80,000 hamburgers will be purchased.
Based on these data, describe the relevant relationship between the price of a hamburger and the quantity consumers are willing to purchase, using a verbal statement, a numerical table, and a graph. Which model do you prefer and why?
what is the answer of this question
1
Expert's answer
2017-03-13T11:40:06-0400
If the price of a hamburger is $4, 20,000 will be bought, at a price of $3, 40,000 hamburgers will be bought, at $2, 60,000 hamburgers will be bought, and at $1, 80,000 hamburgers will be purchased, then the relevant relationship between the price of a hamburger and the quantity consumers are willing to purchase, is:
(Q - 20,000)/(80,000 - 20,000) = (P - 4)/(1 - 4),
(Q - 20,000)/60,000 = (4 - P)/3,
20,000*(4 - P) = Q - 20,000,
80,000 - 20,000P = Q - 20,000,
Qd = 100,000 - 20,000P.
Graph is most preferable model then other, because it clearly shows the relationship between the price and quantity.